# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

### v4

@@ 3.2 Converging properties of MAGPIE and weight selection
$\bm{e}_{H,k}$ $=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{z}_{H,k}^{\prime}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right)$ (18)
$=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
The proof of this theorem appears in Appendix G.1. First, we state Proposition G.1 showing a specific form under which the coarse-grid update is a descent direction. Then we show that the coarse-grid update in Eqn. (17) indeed falls into that form.
Whenever a model is represented on both fine and coarse grids, parameters fixed on the fine grid must have consistent counterparts on the coarse grid. We show below that, with $\bm{W}_{\bm{z}},\bm{W}_{\mathcal{R}},\bm{W}_{\bm{u}}^{H}$ designed as in Eqn. (16), the regularization term can be transferred automatically between levels, requiring no manual tuning for the coarse-grid regularization term $\bm{u}_{H}$.
@@
$\bm{u}_{H}=\bm{W}_{\bm{u}}^{H}\bm{I}_{h}^{H}\bm{u}\,,$
the induced coarse-grid regularization term $\bm{u}_{H}$ ensures that, for each scanning region $k$, there exist mappings
@@
  • The mappings $\bm{e}_{k}(\bm{g})$ and $\bm{e}_{H,k}(\bm{g})$ yield the fine-grid update at $\bm{z}_{k}^{\prime}$ and the coarse-grid update at $\bm{z}_{H,k}^{\prime}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}_{k}^{\prime})$, respectively, for the fine-grid gradient $\bm{g}=\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}(\bm{z}_{k}^{\prime};\,\bm{z}_{k}^{(j)})$.
  • For any coordinate index $r=1,\dots,(m/2)^{2}$, the coarse-grid update satisfies
@@
    Thus, provided the fine-grid update is stable under a properly tuned regularization term, the maximum magnitude of the coarse-grid update is automatically controlled.
The proof of this proposition appears in Appendix H. The key step uses Eqn. (17), which links the coarse-grid update to the fine-grid surrogate gradient.
@@ 4 Numerical Examples
In this section, we present numerical examples to demonstrate the performance of MAGPIE. We compare our solver with classical methods such as rPIE [11] and L-BFGS [28]. For L-BFGS, we set the history size to $5$, which determines the number of recent updates used to approximate the inverse Hessian. To ensure a fair comparison with rPIE, we adopt its regularization term as the fine-grid regularization for MAGPIE, i.e.,
$\bm{u}=\bm{u}^{\texttt{rPIE}}=\alpha\left(\|\bm{Q}\|_{\infty}^{2}-|\bm{Q}|^{2}\right)\,.$
We observe that the computational bottleneck at each iteration stems from the Fourier transforms required to evaluate $\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)$. Since this term is computed only once per scanning region at each iteration, all three algorithms share comparable per-iteration computational costs.
@@
We first describe the problem setup, and then in Sections 4.1, 4.2, and 4.3, we evaluate the performance of the algorithms from three perspectives: the number of multigrid levels, robustness to noise, and the effect of the overlap ratio. Finally, Section 4.4 evaluates the algorithm on a simulated integrated-circuit object to illustrate its practical imaging performance.
Probe:
@@
In our numerical experiments, we use a simulated Fresnel zone plate as the probe (see [29]). The standard probe size is $m=128$. Figure 3 shows the magnitude and phase of the probe. The magnitude displays the characteristic circular structure of a zone plate, while the phase reveals the corresponding alternating concentric patterns. It is worth noting that a large portion of the probe (beyond the center) has a very small magnitude, leading to possible ill-conditioning.
@@
Figure 3: Magnitude (left) and phase (right) of the simulated Fresnel zone plate used as the probe in numerical experiments.
Object:
We test two types of objects. The synthetic object used in our numerical experiments is constructed by combining two standard test images: the Baboon image for magnitude and the Cameraman image for phase. Specifically, the Baboon image is normalized to the range $[0,1]$, representing the object’s magnitude, while the Cameraman image is scaled to span the range $[0,\pi/2]$, providing the object’s phase distribution. The resulting complex-valued object, obtained by multiplying the magnitude by the complex exponential of the phase, exhibits structured patterns and realistic variability in both magnitude and phase. The objects are presented with the resolution of $n=512$. Figure 4 illustrates the magnitude and phase of this synthetic test object.
Figure 4: Magnitude (left) and phase (right) of the synthetic complex-valued object used in numerical experiments. The magnitude is derived from the Baboon image, while the phase is derived from the Cameraman image.
Another synthetic object we use to test our algorithm is more realistic, as shown in Figure 5, which illustrates the magnitude and phase of this test object. This example is inspired by applications in integrated circuit imaging [5].
Figure 5: Magnitude (left) and phase (right) of the realistic complex-valued object used in numerical experiments.
Overlap ratio:
[…]
Noise level:
@@
where the dimensionless parameter $\eta$ defines the noise level: larger $\eta$ lowers the mean photon count and thus raises the relative noise.
In our experiments, the noise level is set to $\eta=0.05$ by default.
Stopping criterion:
@@ 4.1 Effect of multigrid levels on performance
In this experiment, we investigate how the number of coarse levels impacts the performance of MAGPIE. In particular, with $l=1,\cdots,\log_{2}(m)$ levels, we name the corresponding method MAGPIE_$\_l$. Since $m=128$, the maximum level for MAGPIE is $\log_{2}(m)=7$.
In this case, we set the convergence tolerance $\texttt{tol}=10^{-4}$, and the regularization constant $\alpha=0.01$.
The residuals and errors are presented in Figure 6 on the log-log scale. We note that MAGPIE_$\_l$ outperforms rPIE and L-BFGS, and converges faster with each additional coarse level. Therefore, for subsequent tests, we consider only MAGPIE at its deepest level, $l=\log(m)$.
[…]
@@
[…]
@@ 4.2 Stability under noise
[…]
@@
[…]
@@ 4.3 Overlap ratio
[…]
4.4 Realistic object
We extend the previous experiments to a realistic object whose entries near the boundary are mostly zero. This test demonstrates that the performance gain of MAGPIE is not solely due to improved reconstruction near the boundary, where fewer scanning regions overlap.
[…]
@@ 5 Conclusion
While we have employed a specific downsampling strategy, alternative approaches are possible. For example, one may define
$\bm{Q}_{H}=\bm{M}_{h}^{H}(\bm{Q})\,,\quad\bm{z}_{H,k}=\bm{I}_{h}^{H}\bm{z}_{k}\,,\quad\text{and }\ \bm{W}_{\mathcal{R}}=\frac{\bm{I}_{H}^{h}\bm{Q}_{H}}{\bm{Q}}\,,$
where $\bm{M}_{h}^{H}$ selects the minimum-norm entry in each bin. In such variants, the properties established in Section 3.2 still hold, although their analysis becomes more challenging.
Looking ahead, we will extend the theoretical framework to blind ptychography and integrate a minibatch strategy for GPU-accelerated real-time feedback. Moreover, the multigrid-surrogate paradigm developed here is readily adaptable to other large-scale phase retrieval and coherent diffractive imaging applications.
@@ 6 Acknowledgments
Appendix A Computation for Eqn. (2) and Eqn. (4)
@@
Using the chain rule for the cogradient operator defined in Chapter 4.2 of [22], we have
@@
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$
@@
Appendix B Proof for Proposition 2.1, and Theorem 2.2
@@
$\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}\right)\right\|_{2}^{2}\leq\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,.$
Additionally, equality implies $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ except at the zero entries of $\bm{d}_{k}$ or at the zero entries of $\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)$.
@@
$\left\|\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)-\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)\right)\right)\right\|_{2}^{2}\leq\left\|\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)-\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}^{2}\,.$
The inequality and the additional equality outcomes follow from Lemma B.1, which applies to the inequality element-wise. ∎
B.1 Proof for Proposition 2.1
@@
For the second statement, it follows from Proposition B.2 that $\Phi_{k}\left(\bm{z}_{k}\right)\leq\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ for each scanning region. Hence,
@@
    $\bm{z}^{(k+1)}=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,,$
@@
    $\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\leq\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}=\Phi\left(\bm{z}^{(j)}\right)\,.$ (19)
    Then, by Proposition B.2, we have
    $\Phi\left(\bm{z}^{(j+1)}\right)=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)\right\|_{2}^{2}\leq\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,.$ (20)
@@
    Optimality: By Proposition B.2, the inequality Eqn. (20) achieves equality only when $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ except at the zero entries of $\bm{d}_{k}$ or at the zero entries of $\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)$ for all $k=1,2,\dots,N$.
    Since at the zero entries $\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)$, we impose $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ in the Eqn. (8), we have, for all $k=1,2,\dots,N$,
    $\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)$ $=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)\right)\right)\,,$
    $=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right)\right)\,,$
@@
    $\bm{z}^{(j+1)}$ $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,,$
    $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)\right\|_{2}^{2}\,,$
@@
    $\nabla_{\bm{z}}\widetilde{\Phi}\left(\bm{z}^{(j+1)};\bm{z}^{(j+1)}\right)=\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}^{(j+1)}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j+1)}\right)\right)=0\,.$
@@
Appendix C Proof for Theorem 2.3

### v6

@@ 4.2 Converging properties of MAGPIE and weight selection
$\bm{e}_{H,k}=-|\bm{Q}_{H}|^{2}\oslash\left(\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)\oslash\left(\bm{u}_{H}+|\bm{Q}_{H}|^{2}\right)\odot\left(\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$ (18)
The proof of this theorem appears in Appendix H.1. First, we state Proposition H.1 showing a specific form under which the coarse-grid update is a descent direction. Then we show that the coarse-grid update in Eqn. (17) indeed falls into that form. Whenever a model is represented on both fine and coarse grids, parameters fixed on the fine grid must have consistent counterparts on the coarse grid. We show below that, with $\bm{W}_{\bm{z}},\bm{W}_{\mathcal{R}},\bm{W}_{\bm{u}}^{H}$ designed as in Eqn. (16), the regularization term can be transferred automatically between levels, requiring no manual tuning for the coarse-grid regularization term $\bm{u}_{H}$.
@@
$\bm{u}_{H}=\bm{W}_{\bm{u}}^{H}\odot\bm{I}_{h}^{H}\bm{u}\,$,
the induced coarse-grid regularization term $\bm{u}_{H}$ ensures that, for each scanning region $k$, there exist mappings (with a slight abuse of notation)
@@
  • The mappings $\bm{e}_{k}(\bm{g})$ and $\bm{e}_{H,k}(\bm{g})$ yield the fine-grid update at $\bm{z}_{k}^{\prime}$ and the coarse-grid update at $\bm{z}_{H,k}^{\prime}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\odot\bm{z}_{k}^{\prime})$, respectively, for the fine-grid gradient $\bm{g}=\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}(\bm{z}_{k}^{\prime};\,\bm{z}_{k}^{(j)})$.
  • For any index $r=1,\dots,(m/2)^{2}$, the coarse-grid update satisfies
@@
    Thus, provided the fine‐grid update is stable under a properly tuned regularization term, the maximum magnitude of the coarse‐grid update is automatically controlled.
The proof of this proposition appears in Appendix I. The key step uses Eqn. (17), which links the coarse‐grid update to the fine‐grid surrogate gradient.
@@ 5 Numerical Examples
In this section, we present numerical examples to demonstrate the performance of MAGPIE. We compare our solver with classical methods such as rPIE [rPIE] and L-BFGS [LBFGS]. For L-BFGS, we set the history size to $5$, which determines the number of recent updates used to approximate the inverse Hessian. To ensure a fair comparison with rPIE, we adopt its regularization term as the fine-grid regularization for MAGPIE, i.e.,
$\bm{u}=\bm{u}^{\texttt{rPIE}}=\alpha\left(\|\bm{Q}\|_{\infty}^{2}\bm{J}-|\bm{Q}|^{2}\right)\,.$
We observe that the computational bottleneck at each iteration is the evaluation of the revised exit waves $\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)$, which requires Fourier transforms for each scanning region. This term is computed only once per scanning position per iteration, and MAGPIE introduces only additional element-wise operations and restriction/prolongation between levels, whose cost is dominated by the cost of evaluating $\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)$. Moreover, the deeper the level, the smaller the computational domain becomes; at the deepest level, $\log_{2}(m)$, the update for each scan position reduces to a single scalar. As a consequence, the per-iteration cost of MAGPIE is essentially the same as that of rPIE and L-BFGS, while the multilevel corrections provide improved convergence at negligible additional cost, particularly on the deeper levels. The observed speedup is consistent with the fact that coarse-level updates aggregate information from larger neighborhoods, so large-scale (low-frequency) components of the error are damped more efficiently on deeper levels.
@@
We first describe the problem setup, and then in Sections 5.1, 5.2, and 5.3, we evaluate the performance of the algorithms from three perspectives: the number of multigrid levels, robustness to noise, and the effect of the overlap ratio. Finally, Section 5.4 evaluates the algorithm on a simulated integrated-circuit object to illustrate its practical imaging performance.
Probe
@@
In our numerical experiments, we use a simulated Fresnel zone plate as the probe (see [Ptychography]). The standard probe size is $m=128$. Figure 3 shows the probe’s magnitude and phase. The magnitude exhibits the characteristic circular structure of a zone plate, while the phase displays the corresponding alternating concentric patterns. Notably, much of the probe outside the central region has very low magnitude, which can lead to ill-conditioning.
Object
We test two types of objects. The first, a test synthetic object, is constructed by combining two standard test images: the Baboon image for magnitude and the Cameraman image for phase. The Baboon image is normalized to $[0,1]$ to represent magnitude, and the Cameraman image is scaled to $0,\tfrac{\pi}{2}]$ to provide the phase distribution. The second object is a more realistic synthetic model inspired by integrated‐circuit imaging [integrated_circuit]. Both objects are represented at a resolution of $n=512$ and displayed in Figure 3.
@@
Figure 3: Columns (left to right) show the magnitude (top) and phase (bottom) of the three inputs used in our numerical experiments: (1) the simulated Fresnel zone‐plate probe; (2) the test synthetic object (magnitude from the Baboon image; phase from the Cameraman image); and (3) the realistic synthetic object inspired by integrated‐circuit imaging.
Overlap ratio
In our experiments, the overlap ratio is set to 50% by default, i.e., $\texttt{overlap\_ratio}=0.5$.
Noise level
@@
where the dimensionless parameter $\eta$ defines the noise level: larger $\eta$ lowers the mean photon count and thus raises the relative noise. In our experiments, the noise level is set to $\eta=0.05$ by default.
Stopping criterion
@@ 5.1 Effect of multigrid levels on performance
In this experiment, we investigate how the number of coarse levels impacts the performance of MAGPIE. In particular, with $l=1,\cdots,\log_{2}(m)$ levels, we name the corresponding method MAGPIE_ll. Since $m=128$, the maximum level for MAGPIE is $\log_{2}(m)=7$. In this case, we set the convergence tolerance $\texttt{tol}=10^{-4}$, and the regularization constant $\alpha=0.01$.
The residuals and errors are presented in Figure 4 on the log-log scale. We note that MAGPIE_ll outperforms rPIE and L-BFGS, and converges faster with each additional coarse level. Therefore, for subsequent tests, we consider only MAGPIE at its deepest level, $l=\log(m)$.
Figure 4: Log-log plots of residuals (top) and errors (bottom) for L-BFGS, rPIE, and MAGPIE_ll (levels 1 to $\log(m)$) applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05$, $\texttt{overlap\_ratio}=0.5$, $\alpha=0.01$, and $\texttt{tol}=10^{-4}$.
@@
Figure 5: Reconstructions of L-BFGS, rPIE, and MAGPIE (at level $\log(m)$) applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05$, $\texttt{overlap\_ratio}=0.5$, $\alpha=0.01$, and $\texttt{tol}=10^{-4}$.
@@ 5.2 Stability under noise
Figure 6: Log-log plots of residuals for L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05,0.1,0.2$, and $0.4$, $\texttt{overlap\_ratio}=0.5$, $\alpha=0.025$, and $\texttt{tol}=10^{-4}$.
Figure 7: Log-log plots of errors for L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05,0.1,0.2$, and $0.4$, $\texttt{overlap\_ratio}=0.5$, $\alpha=0.025$, and $\texttt{tol}=10^{-4}$.
@@
Figure 8: Reconstructions of L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.4$, $\texttt{overlap\_ratio}=0.5$, $\alpha=0.025$, and $\texttt{tol}=10^{-4}$.
@@ 5.3 Overlap ratio
Another important factor that affects the performance of MAGPIE is the overlapping ratio between adjacent probes, and we are to compare MAGPIE with classical solvers (rPIE and L-BFGS) in various non-default settings of overlapping ratios. The two cases that we study are: $\texttt{overlap\_ratio}=0.25$ (we set $\alpha=0.008$, and $\texttt{tol}=10^{-5}$) and $\texttt{overlap\_ratio}=0.75$ (we set $\alpha=0.04$, and $\texttt{tol}=10^{-4}$). The residuals and errors for both cases are presented in Figure 9. Reconstructions for $\texttt{overlap\_ratio}=0.25$ are showcased in Figure 10, and for $0.75$ are showcased in Figure 11. MAGPIE outperforms rPIE and L-BFGS in all cases.
Figure 9: Log-log plots of residuals and errors for L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe size of $m=128$, using noise level $\eta=0.05$. The first column corresponds to the test with $\texttt{overlap\_ratio}=0.25$, $\alpha=0.008$, and $\texttt{tol}=10^{-5}$, while the second column corresponds to the test with $\texttt{overlap\_ratio}=0.75$, $\alpha=0.04$, and $\texttt{tol}=10^{-4}$.
Figure 10: Reconstructions of L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05$, $\texttt{overlap\_ratio}=0.25$, $\alpha=0.008$, and $\texttt{tol}=10^{-5}$.
Figure 11: Reconstructions of L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05$, $\texttt{overlap\_ratio}=0.75$, $\alpha=0.04$, and $\texttt{tol}=10^{-4}$.
5.4 Realistic synthetic object
We extend the previous experiments to a realistic synthetic object whose entries near the boundary are mostly zero. This test demonstrates that the performance gain of MAGPIE is not solely due to improved reconstruction near the boundary, where fewer scanning regions overlap.
The two cases that we study are: $\texttt{overlap\_ratio}=0.50$ (we set $\alpha=0.03$, and $\texttt{tol}=3\times 10^{-5}$) and $\texttt{overlap\_ratio}=0.75$ (we set $\alpha=0.06$, and $\texttt{tol}=6\times 10^{-5}$).
The residuals and errors for both cases are presented in Figure 12. Reconstructions for $\texttt{overlap\_ratio}=0.50$ are showcased in Figure 13, and for $0.75$ are showcased in Figure 14. MAGPIE outperforms rPIE and L-BFGS in all cases.
Figure 12: Log-log plots of residuals and errors for L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe size of $m=128$, using noise level $\eta=0.05$. The first column corresponds to the test with $\texttt{overlap\_ratio}=0.50$, $\alpha=0.03$, and $\texttt{tol}=3\times 10^{-5}$, while the second column corresponds to the test with $\texttt{overlap\_ratio}=0.75$, $\alpha=0.06$, and $\texttt{tol}=6\times 10^{-5}$.
Figure 13: Reconstructions of L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05$, $\texttt{overlap\_ratio}=0.50$, $\alpha=0.03$, and $\texttt{tol}=3\times 10^{-5}$.
Figure 14: Reconstructions of L-BFGS, rPIE, and MAGPIE applied to a synthetic object ($n=512$) with a probe ($m=128$) using noise level $\eta=0.05$, $\texttt{overlap\_ratio}=0.75$, $\alpha=0.06$, and $\texttt{tol}=6\times 10^{-5}$.
@@ 6 Conclusion
Looking ahead, we will extend the theoretical framework to blind ptychography and integrate a minibatch strategy for GPU-accelerated real‐time feedback. Moreover, the multigrid‐surrogate paradigm developed here is readily adaptable to other large‐scale phase retrieval and coherent diffractive imaging applications.
@@ 7 Acknowledgments
Appendix A Notational convention
For the sake of readability in the computational details, we adopt a slightly informal convention for element-wise division that differs from the main text. Whenever the numerator and denominator have the same dimensions (e.g., vectors or matrices in $\mathbb{C}^{n}$), we write
$\frac{\bm{x}}{\bm{y}}$
to denote element-wise division, that is,
$\frac{\bm{x}}{\bm{y}}\vcentcolon=(x_{i}/y_{i})_{i}.$
[…]
Appendix B Computation for Eqn. (2) and Eqn. (4)
@@ Appendix A Notational convention
Using the chain rule for the cogradient operator defined in Chapter 4.2 of [CRcalculus], we have
@@
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}^{\top}\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$
@@
Appendix C Proof for Proposition 3.1, and Proposition 3.2
@@
$\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}\right)\right\|_{2}^{2}\leq\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,.$
Additionally, equality implies $\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)$ except at the zero entries of $\bm{d}_{k}$ or at the zero entries of $\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}\right)$.
@@
$\left\|\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}\right)-\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}\right)\right)\right)\right\|_{2}^{2}\leq\left\|\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}\right)-\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}^{2}\,.$
The inequality and the additional equality outcomes follow from Lemma C.1, which applies to the inequality element-wise. ∎
C.1 Proof for Proposition 3.1
@@
For the second statement, it follows from Proposition C.2 that $\Phi_{k}\left(\bm{z}_{k}\right)\leq\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ for each scanning region. Hence,
@@
    $\bm{z}^{(k+1)}=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,$,
@@
    $\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\leq\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{(j)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}=\Phi\left(\bm{z}^{(j)}\right)\,.$ (19)
    Then, by Proposition C.2, we have
    $\Phi\left(\bm{z}^{(j+1)}\right)=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)\right\|_{2}^{2}\leq\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,.$ (20)
@@
    Optimality: By Proposition C.2, the inequality Eqn. (20) achieves equality only when $\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)$ except at the zero entries of $\bm{d}_{k}$ or at the zero entries of $\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j+1)}\right)$ for all $k=1,2,\dots,N$.
    Since at the zero entries $\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j+1)}\right)$, we impose $\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)$ in the Eqn. (8), we have, for all $k=1,2,\dots,N$,
    $\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)$ $=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j+1)}\right)\right)\right)\right)\,$,
    $=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)\right)\right)\,$,
@@
    $\bm{z}^{(j+1)}$ $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,$,
    $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)\right\|_{2}^{2}\,$,
@@
    $\nabla_{\bm{z}}\widetilde{\Phi}\left(\bm{z}^{(j+1)};\bm{z}^{(j+1)}\right)=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}^{(j+1)}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j+1)}\right)\right)=0\,.$
@@
Appendix D Proof for Theorem 3.3